# Attention as an Adaptive Filter

### v1

@@ Abstract
We introduce Adaptive Filter Attention (AFA), a novel attention mechanism that incorporates a learnable dynamics model directly into the computation of attention weights. Rather than comparing queries and keys directly, we model the input sequence as discrete observations of a linear stochastic differential equation (SDE). By imposing a linear dynamics model with simultaneously diagonalizable state matrices and noise covariances, we can make use of a closed-form solution to the differential Lyapunov equation to efficiently propagate pairwise uncertainties through the dynamics. Attention naturally arises as the maximum likelihood solution for this linear SDE, with attention weights corresponding to robust residual-based reweightings of the propagated pairwise precisions. Imposing an additional constraint on the state matrix’s eigenvalues leads to a simplified variant with the same computational and memory complexity as standard attention. In the limit of vanishing dynamics and process noise, and using a small-angle approximation, we recover ordinary dot-product attention.
@@ 1 Introduction
Self-attention has become the dominant paradigm for sequence modeling due to its parallelism and scalability. However, this parallelism comes at the cost of losing the recursive structure that RNNs naturally impose, which provides an implicit temporal regularization that encourages consistency over time. State space models (SSMs) offer a complementary approach: they retain temporal structure through shared linear dynamics while supporting parallel training via convolutional implementations. However, most SSMs assume deterministic dynamics, limiting their ability to capture stochasticity and uncertainty.
Our goal is to bridge these two paradigms. We view self-attention as a form of filtering: it selectively averages across inputs, denoising by pooling information from similar states. However, unlike Kalman Filtering, standard attention does not explicitly propagate states through time. This motivates incorporating explicit temporal structure—through known or learned dynamics—into attention, restoring the inductive bias that standard attention loses by forgoing recurrence.
@@
We introduce the model setup and review standard results for propagating covariances through linear SDEs. We then discuss the batch maximum likelihood estimator and its connection to the Kalman Filter. Next, we derive a closed-form solution of the differential Lyapunov equation for diagonalizable systems with simultaneously diagonalizable process and measurement noise covariances. Finally, we discuss an adaptively reweighted MLE that reweights propagated covariances using the pair-wise residuals, and its relationship to the iteratively reweighted least squares (IRLS) algorithm and a robustified form of the Kalman Filter.
@@
We generalize the robustified MLE to a tensor form of attention, and briefly analyze its time and memory complexity. We show that the algorithm can be simplified to a form with the same time and memory complexity as standard attention, through the use of convolutional kernels for the (diagonalized) matrix exponential and pair-wise propagated covariances, and by imposing additional structure on the linear SDE. We show that this formulation reduces to a complex-valued variant of standard attention in the case of a skew-symmetric state matrix and normalized inputs. We then describe an implementation of this algorithm, and an unrolled, recursive approximation for inference. We describe basic experiments on simple simulated linear dynamical systems.
@@
We generalize the SDE model from Section 3, which assumed isotropic noise, to allow for separate noise covariances in the tangential and radial directions. We show that the propagated covariances in this case can be written in closed-form for the case of diagonal plus rank-1 covariances that co-rotate with the dynamics. We derive a form of attention that solves the MLE while respecting the geometric constraints imposed by the assumed dynamics model, and compare it to the standard Transformer architecture.
@@
Appendix A: Background
We briefly overview basic background on linear systems, stochastic processes, the Kalman Filter, adaptive filtering, the attention mechanism, and neural state space models.
@@
We cover some additional topics, including a separable approximation of the propagated precision kernel, discussion of normal plus rank-1 noise covariance, and the expressivity of our structured SDE model.
@@ 2 Preliminaries
We review standard results on vector SDEs and linear Gaussian filtering. The reader familiar with these topics may skip to Section 3. Additional background material can be found in Appendix A.
2.1 Model Setup and Definitions
@@
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)dt+\mathbf{B}d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
where $\mathbf{x}(t)\in\mathbb{R}^{d_{x}}$ is the latent state, $\mathbf{z}_{k}\in\mathbb{R}^{d_{z}}$ is the $k$th measurement, $\mathbf{A}$ is the state matrix, $\mathbf{B}$ is the input matrix, and $\mathbf{C}$ is the output matrix, $d\mathbf{w}(t)\sim\mathcal{N}(\mathbf{0},\mathbf{I}dt)$ is a standard Wiener process (Brownian motion), $\mathbf{Q}:=\mathbf{B}\mathbf{B}^{\top}$ is the continuous-time process noise covariance, and $\mathbf{v}(t_{k})$ is zero-mean Gaussian measurement noise with covariance $\mathbf{R}$, i.e. $\mathbf{v}(t_{k})\sim\mathcal{N}(\mathbf{0},\mathbf{R})$.
With initial condition $\mathbf{x}(t_{i})$, the state at a later time $t_{j}>t_{i}$ can be written explicitly using the matrix exponential and an Itô integral:
$\mathbf{x}(t_{j})=e^{\mathbf{A}(t_{j}-t_{i})}\mathbf{x}(t_{i})+\int_{t_{i}}^{t_{j}}e^{\mathbf{A}(t_{j}-s)}\mathbf{B}d\mathbf{w}(s),$
Letting $\tau=s-t_{i}$, this is:
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}\mathbf{B}d\mathbf{w}(\tau),$
where $\Delta t_{ij}=t_{j}-t_{i}$.
@@
Likewise, if we measure $\mathbf{z}(t_{j})$, and the measurement matrix $\mathbf{C}$ is square and invertible, we can estimate the latent state $\mathbf{x}(t_{i})$, or equivalently $\mathbf{z}^{C}(t_{i}):=\mathbf{C}\mathbf{x}(t_{i})$, by computing a “pulled-back” or "pulled-forward" measurement:
$\mathbf{\hat{z}}_{i}(t_{j}):=\mathbf{C}e^{-\mathbf{A}\Delta t_{ij}}\mathbf{C}^{-1}\mathbf{z}(t_{j}).$
2.2 Propagation of Uncertainty
Recall that the state at time $t_{j}$ can be written as:
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}\,d\mathbf{w}(\tau)\right).$
The covariance of the integral term is:
$\mathbf{V}_{B}(\Delta t_{ij}):=\mathbb{E}\Bigg{[}\left(\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau_{1}}\mathbf{B}\,d\mathbf{w}(\tau_{1})\right)\left(\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau_{2}}\mathbf{B}\,d\mathbf{w}(\tau_{2})\right)^{\top}\Bigg{]}$
$\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{Q}e^{-\mathbf{A}^{\top}\tau}\,d\tau,\quad\mathbf{Q}=\mathbf{B}\mathbf{B}^{\top}$
where the forward-propagated state covariance is:
@@ 2.2 Propagation of Uncertainty
with initial condition $\mathbf{V}_{F}(0)=\mathbf{0}$.
Recall that the estimate of the state is:
$\mathbf{\hat{x}}_{i}(t_{j}):=e^{\mathbf{-A}\Delta t_{ij}}\mathbf{x}(t_{j})=\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau).$
$\mathbf{x}(t_{i})=e^{\mathbf{A}(t_{i}-t_{j})}\mathbf{x}(t_{j})+\int_{t_{j}}^{t_{i}}e^{\mathbf{A}(t_{i}-\tau)}\mathbf{B}d\mathbf{w}(\tau)$
$e^{-\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{j})+\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)\right).$
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i})-\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)$
So:
We again form an estimate:
$\mathbf{\hat{x}}_{i}(t_{j})=e^{-\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{j})=\mathbf{x}(t_{i})-e^{-\mathbf{A}\Delta t_{ij}}\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)$
[…]
Recall that the measurement at time $t_{j}$ is:
$\mathbf{z}(t_{j})=\mathbf{C}e^{\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)\right)+\mathbf{v}(t_{j}).$
The measurement process is distributed as:
$\mathbf{\hat{z}}_{i}(t_{j}):=\mathbf{M}(\Delta t_{ij})\mathbf{z}(t_{j})\quad\text{where}\quad\mathbf{M}(\Delta t):=\mathbf{C}e^{\mathbf{-A}\Delta t}\mathbf{C}^{-1}.$
In terms of the "pulled-forward" state estimate, this is:
$\mathbf{C}\mathbf{\hat{x}}_{i}(t_{j})+\mathbf{M}(\Delta t_{ij})\mathbf{v}(t_{j})$
Hence,
@@
where the propagated measurement covariance is:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\top}+\mathbf{M}(\Delta t_{ij})\mathbf{R}\mathbf{M}^{\top}(\Delta t_{ij})$
2.3 Batch Maximum Likelihood Trajectory Estimation
Taking the derivative with respect to $\mathbf{x}_{i}$:
For the causal case:
@@ 2.3 Batch Maximum Likelihood Trajectory Estimation
$\mathbf{z}(t_{i+1})=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{z}^{C}(t_{i})+\mathbf{C}\int_{0}^{\Delta t}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)+\mathbf{v}(t_{i+1}),$
$\mathbf{z}^{\mathrm{pred}}(t_{i+1}):=\mathbb{E}[\mathbf{z}(t_{i+1})\mid\mathbf{\bar{z}}_{i}]=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{\bar{z}}_{i}.$
2.4 Recursive (Kalman Filter) Formulation
We can express the batch MLE for a trajectory in a recursive (information-form) manner, while accounting for the dynamics of the system. Consider estimating the state $\mathbf{z}^{C}(t_{i})$ given all measurements $\{\mathbf{z}(t_{j})\}_{j\leq i}$.
Prediction (Propagate Prior Through Dynamics).
Suppose we have already incorporated measurements up to time $t_{i-1}$. Let the posterior at $t_{i-1}$ be $\mathbf{\bar{z}}_{i-1}^{+}$ with covariance $\bm{\Lambda}_{\mathbf{V},i-1}^{+}$ and precision $\bm{\Lambda}_{\mathbf{P},i-1}^{+}=(\bm{\Lambda}_{\mathbf{V},i-1}^{+})^{-1}$. Propagating through the linear dynamics:
$\mathbf{z}_{i}=\mathbf{A}\mathbf{z}_{i-1}+\mathbf{q}_{i-1},\quad\mathbf{q}_{i-1}\sim\mathcal{N}(0,\mathbf{Q}),$
the prior for time $t_{i}$ becomes:
$\mathbf{\bar{z}}_{i}^{-}=\mathbf{A}\mathbf{\bar{z}}_{i-1}^{+},\quad\bm{\Lambda}_{\mathbf{V},i}^{-}=\mathbf{A}\bm{\Lambda}_{\mathbf{V},i-1}^{+}\mathbf{A}^{\top}+\mathbf{Q},\quad\bm{\Lambda}_{\mathbf{P},i}^{-}=(\bm{\Lambda}_{\mathbf{V},i}^{-})^{-1},\quad\bm{\eta}_{i}^{-}=\bm{\Lambda}_{\mathbf{P},i}^{-}\,\mathbf{\bar{z}}_{i}^{-}$
Update (Incorporate Measurement).
Now include the measurement at $t_{i}$ with measurement model:
$\mathbf{z}_{i}=\mathbf{C}\mathbf{z}^{-}_{i}+\mathbf{r}_{i},\quad\mathbf{r}_{i}\sim\mathcal{N}(0,\mathbf{R}),$
and define the information-form measurement contribution:
$\mathbf{P}^{C}_{ii}=\mathbf{C}^{\top}\mathbf{R}^{-1}\mathbf{C},\qquad\mathbf{b}_{i}=\mathbf{C}^{\top}\mathbf{R}^{-1}\mathbf{z}_{i}.$
The posterior precision and information vector are then:
$\bm{\Lambda}_{\mathbf{P},i}^{+}=\bm{\Lambda}_{\mathbf{P},i}^{-}+\mathbf{P}^{C}_{ii},\qquad\bm{\eta}_{i}^{+}=\bm{\eta}_{i}^{-}+\mathbf{b}_{i},$
and the posterior mean (MAP estimate) is:
$\mathbf{\bar{z}}_{i}^{+}=(\bm{\Lambda}_{\mathbf{P},i}^{+})^{-1}\bm{\eta}_{i}^{+}.$
Equivalently, in the standard Kalman Filter (covariance) form:
$\mathbf{K}_{i}=\bm{\Lambda}_{\mathbf{V},i}^{-}\mathbf{C}^{\top}(\mathbf{C}\bm{\Lambda}_{\mathbf{V},i}^{-}\mathbf{C}^{\top}+\mathbf{R})^{-1},\quad\mathbf{\bar{z}}_{i}^{+}=\mathbf{\bar{z}}_{i}^{-}+\mathbf{K}_{i}(\mathbf{z}_{i}-\mathbf{C}\mathbf{\bar{z}}_{i}^{-}),\quad\bm{\Lambda}_{\mathbf{V},i}^{+}=(\mathbf{I}-\mathbf{K}_{i}\mathbf{C})\bm{\Lambda}_{\mathbf{V},i}^{-}.$
2.5 Closed-form Precision Matrix Calculation for Diagonalizable Matrices
Note that $(e^{-\bm{\Lambda}t}\mathbf{G}e^{-\bm{\Lambda}^{\dagger}t})_{kl}=\mathbf{G}_{kl}e^{-\lambda_{k}t}e^{-\bar{\lambda}_{l}t}=\mathbf{G}_{kl}e^{-(\lambda_{k}+\bar{\lambda}_{l})t}$. Hence, the $kl$th component of $\bm{\Lambda}_{\mathbf{V}B}(\Delta t)$ is:
$\bm{\Lambda}_{\mathbf{V}B}(\Delta t)_{kl}=\mathbf{G}_{kl}\int_{0}^{|\Delta t|}e^{-(\lambda_{k}+\bar{\lambda}_{l})\tau}d\tau$
$\begin{cases}\mathbf{G}_{kl}\,\dfrac{1-e^{-(\lambda_{k}+\bar{\lambda}_{l})|\Delta t|}}{\lambda_{k}+\bar{\lambda}_{l}},&\lambda_{k}+\bar{\lambda}_{l}\neq 0\\[5.69054pt] \mathbf{G}_{kl}\,|\Delta t|,&\lambda_{k}+\bar{\lambda}_{l}=0\end{cases}$
For the causal case, the forward-propagated covariance is:
$\mathbf{S}\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)\mathbf{S}^{\dagger}$
where
@@ 2.5 Closed-form Precision Matrix Calculation for Diagonalizable Matrices
$\mathbf{V}_{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\dagger}+\mathbf{M}(\Delta t_{ij})\mathbf{R}\mathbf{M}^{\dagger}(\Delta t_{ij}),\quad\text{where}\quad\mathbf{M}(\Delta t_{ij})=\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}.$
If $\mathbf{C}=\mathbf{S}\bm{\Lambda}_{C}\mathbf{S}^{-1}$, this becomes: $\mathbf{M}(\Delta t_{ij})=\mathbf{S}\bm{\Lambda}_{C}e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{C}^{-1}\mathbf{S}=\mathbf{S}e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{S}^{-1}$. If, in addition, $\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{\dagger}$, then:
$\mathbf{S}\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})\mathbf{S}^{\dagger}$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\bm{\Lambda}_{R}$
(Stable dynamics attenuate the effect of measurement noise.)
@@
For the non-causal case,
So,
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\text{sign}(\Delta t_{ij})\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{Q}\mathcal{K}(\bm{\Lambda},\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R},$
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\sigma_{V}^{2}(\Delta t_{ij})\mathbf{I},\quad\text{where}\quad\sigma_{V}^{2}(\Delta t_{ij}):=\text{sign}(\Delta t_{ij})\lambda_{C}^{2}\sigma^{2}\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}+e^{-2\alpha\Delta t_{ij}}\eta^{2}$
$\dot{\lambda}_{vk}(t)=2\mathrm{Re}(\lambda_{i})\lambda_{vk}(t)+\lambda_{\Omega k}-\frac{\lambda_{Ck}^{2}}{\lambda_{\Gamma k}}\lambda_{vk}(t)^{2}$
Letting $\alpha_{k}=\mathrm{Re}(\lambda_{k})$, $\beta_{k}=\sqrt{\alpha_{k}^{2}+\frac{\lambda_{Qk}}{\lambda_{Rk}}\lambda_{Ck}^{2}}$, $\gamma_{k}=\frac{1}{\beta_{k}}\log\left(\frac{\beta_{k}-\alpha_{k}}{\beta_{k}+\alpha_{k}}\right)$, the $k$th element of $\bm{\Lambda}_{\mathbf{V}}$ becomes:
$\lambda_{vk}(\Delta t_{ij})=\frac{\lambda_{Rk}}{\lambda_{Ck}^{2}}\left(\beta_{k}\tanh\left(\frac{\beta_{k}}{2}\left(\Delta t_{ij}-\gamma_{k}\right)\right)+\alpha_{k}\right)$

### v2

@@ Abstract
We introduce Adaptive Filter Attention (AFA), a novel attention mechanism that incorporates a learnable dynamics model directly into the computation of attention weights. Rather than comparing queries and keys directly, we model the input sequence as discrete observations of a linear stochastic differential equation (SDE). By imposing a linear dynamics model with simultaneously diagonalizable state matrices and noise covariances, we can make use of a closed-form solution of the differential Lyapunov equation to efficiently propagate pairwise uncertainties through the dynamics. Attention naturally arises as the maximum likelihood solution for this linear SDE, with attention weights corresponding to robust residual-based reweightings of the propagated pairwise precisions. To obtain a simplified variant with the same computational and memory complexity as standard attention, we further constrain the state matrix’s eigenvalues by tying their real parts to a shared scalar, i.e. enforcing isotropic decay. In the limit of vanishing dynamics and process noise, and using a small-angle approximation, we recover a complex-valued form of ordinary dot-product attention with rotary positional encodings.
@@ 1 Introduction
Self-attention has become the dominant paradigm for sequence modeling due to its parallelism and scalability. However, this parallelism comes at the cost of losing the recursive structure that RNNs naturally impose, which provides an implicit temporal regularization that encourages consistency over time. State space models (SSMs) offer a complementary approach: they retain temporal structure through shared linear dynamics while supporting parallel training via convolutional implementations. However, SSMs generally assume deterministic dynamics, limiting their ability to capture stochasticity and uncertainty.
Our goal is to bridge these two paradigms. We view self-attention as a form of filtering: it selectively averages across inputs, denoising by pooling information from similar states. However, unlike Kalman Filtering, standard attention does not explicitly propagate states through time. This motivates incorporating explicit temporal structure—through known or learned dynamics—into attention, restoring the inductive bias that standard attention loses by forgoing recurrence. Unlike SSMs, which are typically formulated as deterministic ODEs with inputs entering explicitly through control terms, we instead treat the entire dataset as arising from a stochastic differential equation (SDE), i.e. an autonomous system driven by noise.
@@
We introduce the model setup and review standard results for propagating covariances through linear SDEs. We then discuss the batch maximum likelihood estimator and the closed-form solution of the differential Lyapunov equation for diagonalizable systems with simultaneously diagonalizable process and measurement noise covariances. Finally, we discuss an adaptively reweighted MLE that reweights propagated covariances using the pair-wise residuals.
@@
We generalize the robustified MLE to a tensor form of attention, and briefly analyze its time and memory complexity. We show that the algorithm can be simplified to a form with the same time and memory complexity as standard attention, through the use of convolutional kernels for the (diagonalized) matrix exponential and pair-wise propagated covariances, and by imposing additional structure on the linear SDE. We show that, in the case of a skew-symmetric state matrix and normalized inputs, this reduces to a complex-valued variant of standard attention with a complex-valued positional encoding equivalent to that in xPos and RetNet. We then describe an implementation of this algorithm, and an unrolled, recursive approximation for inference. We describe basic experiments on simple simulated linear dynamical systems.
@@
We generalize the SDE model from Section 3, which assumed isotropic noise, to allow for separate noise covariances in the tangential and radial directions (which we call a radial-tangential SDE). We show that the propagated covariances in this case can be written in closed-form since the assumed diagonal plus rank-1 covariances co-rotate with the dynamics. This formulation naturally yields a maximum likelihood estimator (MLE) on the hypersphere, and a corresponding attention mechanism. From this perspective, the Transformer can be interpreted geometrically, as an approximate solution of the MLE under the constraints imposed by the radial-tangential SDE model.
@@
Appendix A: Background Additional details for the material in the preliminaries section.
@@
Extensions of AFA, a recurrent formulation, generative modeling, an extension to control, an extension to higher dimensional data, discussion of normal plus rank-1 noise covariance, and discussion of the expressivity of our structured SDE model.
@@ 2 Preliminaries
This section introduces the setup of our filtering problem and key ideas from linear Gaussian filtering on which we base the attention mechanism developed in Section 3. Detailed derivations can be found in Appendix A.
2.1 Model Setup
@@
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)dt+\mathbf{G}d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
with the following definitions:
  • $\mathbf{x}(t)\in\mathbb{R}^{d}$: latent state at time $t$,
  • $\mathbf{z}(t_{k})\in\mathbb{R}^{d}$: $k$th measurement,
  • $\mathbf{A},\mathbf{G},\mathbf{C}\in\mathbb{R}^{d\times d}$: state, input, and output matrices,
  • $d\mathbf{w}(t)\sim\mathcal{N}(\mathbf{0},\mathbf{I}\,dt)$: standard Wiener process (Brownian motion),
  • $\mathbf{Q}:=\mathbf{G}\mathbf{G}^{\top}$: continuous-time process noise covariance,
  • $\mathbf{v}(t_{k})\sim\mathcal{N}(\mathbf{0},\mathbf{R})$: zero-mean Gaussian measurement noise with covariance $\mathbf{R}$.
@@
Likewise, if we measure $\mathbf{z}(t_{j})$, and assuming the measurement matrix $\mathbf{C}$ is square and invertible, we can estimate the latent state $\mathbf{x}(t_{i})$, or equivalently $\mathbf{z}^{C}(t_{i}):=\mathbf{C}\mathbf{x}(t_{i})$, by computing a "pulled-forward" measurement:
$\mathbf{\hat{z}}_{i}(t_{j}):=\mathbf{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{z}(t_{j}),\quad\text{where: }\,\mathbf{\Phi}_{\mathbf{z}}(\Delta t):=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}$
In the special case that $\mathbf{A}$ and $\mathbf{C}$ commute, $\mathbf{\hat{z}}_{i}(t_{j})=\mathbf{\hat{x}}_{i}(t_{j})$.
The "pulled-forward" estimate is distributed around the true state as:
$\mathbf{\hat{x}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{x}(t_{i}),\mathbf{V}(|\Delta t_{ij}|)\right)$
where the forward-propagated covariance is:
$\mathbf{V}(|\Delta t_{ij}|)=\int_{0}^{|\Delta t_{ij}|}e^{\mathbf{A}s}\mathbf{Q}e^{\mathbf{A}^{\top}s}\,ds.$
@@
$\frac{d}{ds}\mathbf{V}(s)=\mathbf{A}\mathbf{V}(s)+\mathbf{V}(s)\mathbf{A}^{\top}+\mathbf{Q},$
The pulled-forward measurement estimate is distributed as:
@@
where the forward-propagated measurement covariance is:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\top}+\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{R}\bm{\Phi}_{\mathbf{z}}^{\top}(\Delta t_{ij})$
We can also define corresponding precision matrices $\mathbf{P}(\Delta t_{ij})=\big(\mathbf{V}(\Delta t_{ij})\big)^{-1}$ and $\mathbf{P}^{C}(\Delta t_{ij})=\big(\mathbf{V}^{C}(\Delta t_{ij})\big)^{-1}$.
2.2 Batch Maximum Likelihood Estimation
We wish to estimate the trajectory of the system from $m$ measurements taken at times $\{t_{1},...,t_{m}\}$, which we denote $\{\mathbf{z}(t_{1}),...,\mathbf{z}(t_{m})\}$. We can do so by solving a separate maximum likelihood estimation (MLE) problem for each $\mathbf{x}(t_{i})$.
To simplify the notation, let $\mathbf{x}_{i}=\mathbf{x}(t_{i})$, $\mathbf{\hat{x}}_{ij}=\mathbf{\hat{x}}_{i}(t_{j})$ and $\mathbf{P}_{ij}=\mathbf{P}(\Delta t_{ij})$, etc. The (causal) MLE solution for $\mathbf{z}^{C}(t_{i})$ is a precision-weighted estimator:
$\mathbf{\bar{z}}_{i}:=\left(\sum_{j\leq i}\mathbf{P}_{ij}^{C}\right)^{-1}\sum_{j\leq i}\mathbf{P}_{ij}^{C}\mathbf{\hat{z}}_{ij},$
@@ 2.2 Batch Maximum Likelihood Estimation
Computing $\mathbf{P}^{C}(\Delta t_{ij})$ is generally expensive, as it involves integration and matrix inversion. However, if we assume that $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1},\mathbf{Q}=\mathbf{S}\bm{\Lambda}_{Q}\mathbf{S}^{\dagger}\in\mathbb{R}^{d\times d}$, where $\bm{\Lambda},\bm{\Lambda}_{Q}\in\mathbb{C}^{d\times d}$ are diagonal matrices and $\mathbf{S}\in\mathbb{C}^{d\times d}$ an invertible matrix, the backward-propagated covariance becomes:
$\mathbf{V}_{F}(|\Delta t_{ij}|)=\mathbf{S}\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)\mathbf{S}^{\dagger}$
where:
@@
(where division by matrices denotes element-wise division).
If, in addition, $\mathbf{C}=\mathbf{S}\bm{\Lambda}_{C}\mathbf{S}^{-1},\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{\dagger}$, then:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{S}\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})\mathbf{S}^{\dagger}$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\bm{\Lambda}_{R}$